# Interpretation of 95 GeV Excess within the Georgi-Machacek Model in Light of Positive Definiteness Constraints

### v1

@@ 5 Numerical results
We would like to identify the parameters that can yield a light CP-even Higgs at 95 GeV to account for the excess signals reported by the ATLAS, CMS, and LEP collaborations. Realistic particle spectra, branching ratios, total widths of the scalars, and current experimental constraints, such as the mass of the SM-like Higgs, the couplings of the Higgs to vector bosons and fermions (cited in ATLAS:kappaWZ ; CMS:kappaWZ ; PDG:2022 ), and B-meson rare decays sensitive to new physics (cited in B-physics ), can be calculated using the GMCALC package GMCal (specifically designed for the Georgi-Machacek model) or the SPheno package SPheno1 ; SPheno2 (which, when combined with SARAH SARAH1 ; SARAH2 ; SARAH3 ; SARAH4 ; SARAH5 , can compute the spectrum of general new physics models). Given the rich phenomenology associated with additional Higgs bosons in the Georgi-Machacek model, we utilize the HiggsBounds-5.10.0 package HiggsBounds1 ; HiggsBounds2 ; HiggsBounds3 ; HiggsBounds4 to ensure that our parameter choices comply with the bounds on additional Higgs bosons at colliders. Furthermore, we employ the HiggsSignals-2.6.2 package HiggsSignals1 ; HiggsSignals2 ; HiggsSignals3 ; HiggsTools to verify that the next-to-lightest CP-even Higgs boson mimics the behavior of the SM-like Higgs. Generally speaking, the input parameters include $\lambda_{1,2,3,4,5}$, cubic coupling parameters with mass dimension $\mu_{1,2}$, and mixing angle $\sin{\theta}$ (for Yukawa couplings), and the parameter $\lambda_{1}$ can be replaced by mixing angle $\alpha$. The range of these free input parameters is chosen as
@@
Figure 1: The signal strength of di-photon and $b\bar{b}$ excesses at 95 GeV. The different colors indicate the maximum values that each input parameter point can reach via RGE. All points in the left panels fulfill the positive definiteness constraints at large field value regions but do not meet the tree-level bounded-from-below bounds at the EW scale. The points in the middle panels are completely opposite to those on the left. However, the panels on the right show points that satisfy both the tree-level bounded-from-below bounds at the EW scale and the positive definiteness constraints at large field value regions. All points in the upper panels can successfully provide a 125 GeV SM-like Higgs boson and satisfy the theoretical perturbative unitarity bounds at the EW scale. The points in the middle three panels further need to satisfy constraints related to B-meson rare decay and the oblique parameters ‘TSU’. Additionally, the lower three figures are the result of considering the constraints related to the Higgs boson, as compared to the middle panels.
We conduct a numerical scan and present our results in Fig. 1. In this figure, each set of input parameters must provide a next-to-lightest CP-even Higgs boson with a mass of 125 GeV, which serves as the SM-like Higgs boson, and a lightest CP-even Higgs boson to successfully explain the excess signals. We require that the points shown in Fig. 1 must survive the theoretical perturbative unitarity bounds at the EW scale. Since the number of points obtained by the Markov Chain Monte Carlo (MCMC) method during the numerical scan is excessively large, we only display the points that can offer a quite good interpretation of the 95 GeV excess signals.
The Fig. 1 shows the sample distributions of di-photon and $b\bar{b}$ excesses at a 95 GeV signal strength provided by our work, where we only show the points that can simultaneously interpret these two excesses in the 2 $\sigma$ significance region. And the different colors indicate the maximum values each input parameter point can reach via RGE. All the panels on the left in Fig. 1 (actually for all the plots in this work) show points that fulfill the positive definiteness constraints at large field value regions, but do not meet the tree-level bounded-from-below bounds at the EW scale. The points in the middle panels are completely opposite to those on the left. The panels on the right, however, show points that satisfy both the tree-level bounded-from-below bounds at the EW scale and the positive definiteness constraints at large field value regions. We observe that many parameter points, previously excluded by tree-level bounded-from-below constraints, can actually satisfy the positive definiteness constraints for the scalar potential when considering the one-loop RG-improved scalar potential at large field values regions. Conversely, a portion of the parameter space that meets the tree-level bounded-from-below constraints (using EW scale coupling inputs) should, in fact, be ruled out by the positive definiteness constraints of the effective potential at large field values. Indeed, this situation between the positive definiteness constraints at large field value regions and the tree-level bounded-from-below constraints is completely consistent with the conclusion of GM-pd .
The sample points shown in the upper panels of Fig. 1 can successfully provide a 125 GeV SM-like Higgs boson and satisfy the theoretical perturbative unitarity bounds at the EW scale, regardless of simultaneously explaining the 95 GeV excesses with a 2 $\sigma$ significance. It is evident that these two 95 GeV excesses can be well explained in the GM model, especially when considering the positive definiteness constraints in large field value regions and relaxing the tree-level bounded-from-below constraints. There is a slight decrease in the surviving parameter space in the middle three panels of Fig. 1 compared to the upper ones, as we incorporate the constraints of B-meson rare decay and the oblique parameters “TSU”. Due to the severe suppression of the interaction between the quark-phobic $SU(2)_{V}$ quintuple scalar and quarks, and the weak interaction between the gauge-phobic $SU(2)_{V}$ triplet scalar and fermions in the so-called “alignment limit” of the GM model, the constraints of B-meson rare decay are not very strict. Moreover, considering the invariance of custodial symmetry at the tree level in the GM model, the constraints of the oblique parameters “TSU” are also quite weak, which is also the benchmark feature of the GM model. Additionally, the lower three images are the result of considering the constraints of extra Higgs bosons searching from collider data checked by the package of HiggsBounds-5.10 and the bounds of the specific properties of the SM-like Higgs boson, which is the heavier singlet scalar $h$ in this work, verified by the package of HiggsSignal-2.6.2. By comparing two sets of upper panels, it becomes evident that the Higgs-related constraint has the most significant impact on the parameter space, particularly for parameter points that solely satisfy the tree-level vacuum stability constraint at the electroweak scale. Although many parameter points that satisfy positive definiteness bounds in the large field value regions have been excluded, the remaining points can still adequately explain the 95 GeV excess signals.
Figure 2: The upper three panels of the figure illustrate the mass distribution of the light Higgs boson and its corresponding signal strength. The lower three panels depict the mass ranges for both the SM-like Higgs boson and the doubly charged Higgs boson, with different colors indicating the VEVs of the triplet fields.
After previous discussions, we can reach the important conclusion that the bounds of positive definiteness at large field value regions should be taken into account in the first place, rather than the constraints of bounded-from-below at the EW scale. To tell that more easily, we will conduct a specific analysis of the parameter space that satisfies the stringent theoretical and experimental constraints and can provide a very good explanation for the 95 GeV excess signals in the following part. Hence, all points in Fig. (2,3) not only successfully account for a 125 GeV SM-like Higgs boson but also satisfy several critical constraints: the theoretical bounds of perturbative unitarity at the EW scale, the constraints related to B-meson rare decays, and the requirements imposed by the oblique parameters ”TSU” as well as those associated with the Higgs boson itself. Similar to Fig. 1, the points in the left panels of Fig.(2,3) fulfill the bounds of positive definiteness at large field value regions but can not satisfy the constraints of bounded-from-below at the EW scale, those in the middle panels are the opposite, and those in the right panels can meet these two constraints. In other words, the points presented in Fig. (2,3) are detailed in the lower three panels of Fig. 1.
@@
The lower three panels of Fig. 2 depict the mass ranges for both the SM-like Higgs boson and the doubly charged Higgs boson, with different colors indicating the VEVs of the triplet fields. For points fulfilling the constraints of positive definiteness in large field value regions (shown in the lower left and lower right panels), the mass of the doubly charged Higgs boson varies from 200 GeV to 340 GeV, whereas for those satisfying the bounded-from-below conditions at the EW scale, it ranges only from 220 GeV to 250 GeV; the lower mass region is consistent with the tendency of $H_{5}$ provided by 95-38 . The mass of the doubly charged Higgs slowly increases with that of the SM-like Higgs and shows a significant positive correlation with the VEVs of the triplets which varies from 3 GeV to 30 GeV. Interestingly, in the lower left panel, there exists an independent parameter space that can provide a wide mass range for the doubly charged Higgs, particularly when the VEVs of the triplets are around 16 GeV.
@@
Figure 3: The upper panels display the specific values of the branching ratios for the common B-meson rare decays in the $B\rightarrow\tau\nu$ channel, the $B\rightarrow X_{S}$ channel, and the $B_{S}^{0}\rightarrow\mu^{+}\mu^{-}$ channel. The values of the effective coupling for the interactions between the SM-like Higgs boson and third-generation quarks, as well as the Z vector boson are presented in the lower panels.
@@
All the points presented in Figs. 2 and 3 are detailed in the lower three panels of Fig. 1. This means that the points in Fig. 3 have taken into account the constraints from B-meson rare decays and the effective coupling of the SM-like Higgs boson. So it is clear that all the parameter space shown in Fig. 3 can survive well under the experimental constraints from the common B-meson rare decays (in the $B\rightarrow\tau\nu$ channel, the $B\rightarrow X_{S}$ channel, and the $B_{S}^{0}\rightarrow\mu^{+}\mu^{-}$ channel) and the effective coupling of the SM-like Higgs boson (the interactions between the SM-like Higgs boson and the top quark $\kappa_{t}$, bottom quark $\kappa_{b}$, and the Z vector boson $\kappa_{ZZ}$).
@@
Thanks to the suppressed interaction of the quark-phobic $SU(2)_{V}$ quintuplet scalar with quarks and the weakly coupled nature of the gauge-phobic $SU(2)_{V}$ triplet scalar with fermions in the “alignment limit” of the GM model, constraints from B-meson rare decays are relatively weak even when we apply the newest data PDG:2024 , as demonstrated in the upper three panels of Fig. 3. Consequently, most of the parameter space can survive the incoming collider experiments, such as the High-Luminosity LHC (HL-LHC) HL-LHC1 ; HL-LHC2 . On the contrary, constraints from the precision measurements of Higgs coupling are very stringent for the GM model, which extends the SM with triplet scalar fields. Although the points in the lower panels of Fig. 3 currently meet all bounds, the parameter space will shrink significantly once future collider experiments are considered. In particular, applying the predicted bounds from the Circular Electron-Positron Collider (CEPC)  CEPC1 ; CEPC2 ; CEPC3 (with 0.01 accuracy for top and bottom quarks and 0.001 accuracy for Z bosons), both the lower middle and right panels would be entirely excluded, while only a minimal portion of the parameter space would remain viable for the lower left panel.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China (NSFC) No. 12447167, by the Joint Fund of Henan Province Science and Technology R&\&&D Program No.225200810092, by the Joint Fund of Henan Province Science and Technology R&\&&D Program No.225200810030, by the Startup Research Fund of Henan Academy of Sciences No.231820011, by the Basic Research Fund of Henan Academy of Sciences No.240620006, by the Graduate Innovation Fund of Henan Academy of Sciences No.243320031.

### v2

@@ 5 Numerical results
We would like to identify the parameters that can yield a light $CP$-even Higgs at 95 GeV to account for the excess signals reported by the ATLAS, CMS, and LEP collaborations. Realistic particle spectra, branching ratios, total widths of the scalars, and current experimental constraints, such as the mass of the SM-like Higgs, the couplings of the Higgs to vector bosons and fermions (cited in ATLAS:kappaWZ ; CMS:kappaWZ ; PDG:2022 ), and B-meson rare decays sensitive to new physics (cited in B-physics ), can be calculated using the GMCALC package GMCal (specifically designed for the Georgi-Machacek model) or the SPheno package SPheno1 ; SPheno2 (which, when combined with SARAH SARAH1 ; SARAH2 ; SARAH3 ; SARAH4 ; SARAH5 , can compute the spectrum of general new physics models). Given the rich phenomenology associated with additional Higgs bosons in the Georgi-Machacek model, we utilize the HiggsBounds-5.10.0 package HiggsTools ; HiggsBounds1 ; HiggsBounds2 ; HiggsBounds3 ; HiggsBounds4 to ensure that our parameter choices comply with the bounds on additional Higgs bosons at colliders. Furthermore, we employ the HiggsSignals-2.6.2 package HiggsSignals1 ; HiggsSignals2 ; HiggsSignals3 ; HiggsTools to verify that the next-to-lightest $CP$-even Higgs boson mimics the behavior of the SM-like Higgs.
Generally speaking, the input parameters include quadratic couplings $\lambda_{1,2,3,4,5}$, cubic coupling parameters with mass dimension $\mu_{1,2}$ and the mixing angle $\sin{\theta}$. Specifically, the input parameters $\mu_{1,2}$ can be alternatively expressed in terms of the singlet masses $m_{h}$ and $m_{H}$ to more easily facilitate the matching with their masses, as detailed in Eq. 48. Besides In order to better control the degree of mixing between the two singlets, we can replace $\lambda_{1}$ with the mixing angle of two singlets $\alpha$ as elucidated by Eqs. (47 and 57).The range of these free input parameters is chosen as
@@
Figure 1: The signal strength of di-photon and $b\bar{b}$ excesses at 95 GeV. The different colors indicate the maximum values that each input parameter point can reach via RGE. All points in the top panels fulfill the positive definiteness constraints at large field value regions but do not meet the tree-level bounded-from-below bounds at the EW scale. The points in the middle two panels are completely opposite to those in the top. However, the panels in lower panels show points that satisfy both the tree-level bounded-from-below bounds at the EW scale and the positive definiteness constraints at large field value regions. All points in the left panels can successfully provide a 125 GeV SM-like Higgs boson and satisfy constraints of the theoretical perturbative unitarity bounds at the EW scale, B-meson rare decay and the oblique parameters ‘TSU’. Additionally, the right three figures are the result of considering the constraints related to the Higgs boson, as compared to the left panels.
@@
The Fig. 1 shows the sample distributions of signal strength for di-photon and $b\bar{b}$ excesses at 95 GeV provided by our work. And the different colors indicate the values of maximum scale that each input parameter point can reach via RGE. The two panels at the top in Fig. 1 (actually for all the upper panels on different plots in this work) show points that fulfill the positive definiteness constraints at large field value regions, but do not meet the tree-level bounded-from-below bounds at the EW scale. The points in the middle two panels are completely opposite to those on the top. The panels on the bottom, however, show points that satisfy both the tree-level bounded-from-below bounds at the EW scale and the positive definiteness constraints at large field value regions. We observe that many parameter points, previously excluded by tree-level bounded-from-below constraints, can actually satisfy the positive definiteness constraints for the scalar potential when considering the one-loop RG-improved scalar potential at large field values regions. Conversely, a portion of the parameter space that meets the tree-level bounded-from-below constraints (using EW scale coupling inputs) should, in fact, be ruled out by the positive definiteness constraints of the effective potential at large field values. Indeed, this situation between the positive definiteness constraints at large field value regions and the tree-level bounded-from-below constraints is completely consistent with the conclusion of GM-pd .
The sample points shown in the left three panels of Fig. 1 can successfully provide a 125 GeV SM-like Higgs boson and satisfy the bounds of theoretical perturbative unitarity, B-meson rare decay and the oblique parameters “TSU” at the EW scale, regardless of simultaneously explaining the 95 GeV excesses in 2 $\sigma$. It is evident that these two 95 GeV excesses can be well explained in the GM model, especially when considering the positive definiteness constraints in large field value regions and relaxing the tree-level bounded-from-below constraints. Due to the severe suppression of the interaction between the quark-phobic $SU(2)_{V}$ quintuple scalar and quarks, and the weak interaction between the gauge-phobic $SU(2)_{V}$ triplet scalar and fermions in the so-called “alignment limit” of the GM model, the constraints of B-meson rare decay are not very strict. Moreover, considering the invariance of custodial symmetry at the tree level in the GM model, the constraints of the oblique parameters “TSU” are also quite weak, which is also the benchmark feature of the GM model. Additionally, the right three panels are the result of considering the constraints of extra Higgs bosons searching from collider data checked by the package of HiggsBounds-5.10 and the bounds of the specific properties of the SM-like Higgs boson, which is the heavier singlet scalar $h$ in this work, verified by the package of HiggsSignal-2.6.2. By comparing set of left panels, it becomes evident that the Higgs-related constraint has the most significant impact on the parameter space, particularly for parameter points that solely satisfy the tree-level vacuum stability constraint at the EW scale. Although many parameter points that satisfy positive definiteness bounds in the large field value regions have been excluded, the remaining points can still adequately explain the 95 GeV excess signals.
Figure 2: The left three panels of the figure illustrate the mass distribution of the light Higgs boson and its corresponding signal strength. The right three panels depict the mass ranges for both the SM-like Higgs boson and the doubly charged Higgs boson, with different colors indicating the VEVs of the triplet fields.
After previous discussions, we can reach the important conclusion that the bounds of positive definiteness at large field value regions should be taken into account in the first place, rather than the constraints of bounded-from-below at the EW scale. To tell that more easily, we will conduct a specific analysis of the parameter space that satisfies the stringent theoretical and experimental constraints and can provide a very good explanation for the 95 GeV excess signals in the following part. Hence, all points in Fig. (2,3) not only successfully account for a 125 GeV SM-like Higgs boson but also satisfy several critical constraints: the theoretical bounds of perturbative unitarity at EW scale, the constraints related to B-meson rare decays, and the requirements imposed by the oblique parameters “TSU” as well as those associated with the Higgs boson itself. More precisely, the points presented in Fig. (2,3) are detailed in the lower three panels of Fig. 1. Similar to Fig. 1, the points in the upper panels of Fig.(2,3) fulfill the bounds of positive definiteness at large field value regions but can not satisfy the constraints of bounded-from-below at the EW scale, those in the middle panels are the opposite, and points in the lower panels can meet these two constraints.
[…]
@@
The right three panels of Fig. 2 depict the mass ranges for both the SM-like Higgs boson and the doubly charged Higgs boson, with different colors indicating the VEVs of the triplet fields. For points fulfilling the constraints of positive definiteness in large field value regions (shown in the upper right and the lower right panels), the mass of the doubly charged Higgs boson varies from 200 GeV to 340 GeV, whereas for those satisfying the bounded-from-below conditions at the EW scale (shown in the middle right panel), it ranges only from 220 GeV to 250 GeV; the narrow mass region is consistent with the tendency of $H_{5}$ provided by 95-38 . The mass of the doubly charged Higgs slowly increases with that of the SM-like Higgs and shows a significant positive correlation with the VEVs of the triplets which varies from 3 GeV to 30 GeV.
@@
Figure 3: The left panels display the specific values of the branching ratios for the common B-meson rare decays in the $B\rightarrow\tau\nu$ channel, the $B\rightarrow{X_{s}\gamma}$ channel, and the $B_{s}^{0}\rightarrow\mu^{+}\mu^{-}$ channel. The values of the effective coupling for the interactions between the SM-like Higgs boson and third-generation quarks, as well as the Z vector boson are presented in the right panels.
@@
All the points in Fig. 3 have taken into account the constraints from B-meson rare decays, oblique parameters “TSU” and the effective couplings of the SM-like Higgs boson. So it is clear that all the parameter space shown in Fig. 3 can survive well under the experimental constraints from the common B-meson rare decays (in the $B\rightarrow\tau\nu$ channel, the $B\rightarrow{X_{s}\gamma}$ channel, and the $B_{s}^{0}\rightarrow\mu^{+}\mu^{-}$ channel) and the effective coupling of the SM-like Higgs boson (the interactions between the SM-like Higgs boson and the top quark $\kappa_{t}$, bottom quark $\kappa_{b}$, and the Z vector boson $\kappa_{ZZ}$).
@@
Thanks to the suppressed interaction of the quark-phobic $SU(2)_{V}$ quintuplet scalar with quarks and the weakly coupled nature of the gauge-phobic $SU(2)_{V}$ triplet scalar with fermions in the “alignment limit” of the GM model, constraints from B-meson rare decays are relatively weak even when we apply the newest data PDG:2024 , as demonstrated in the left three panels of Fig. 3. Consequently, most of the parameter space can survive the incoming collider experiments, such as the High-Luminosity LHC (HL-LHC) HL-LHC1 ; HL-LHC2 .
On the contrary, constraints from the precision measurements of Higgs coupling are very stringent for the GM model, which extends the SM with triplet scalar fields. Although the points in the right three panels of Fig. 3 currently meet all bounds, the parameter space will shrink significantly once future collider experiments are considered. In particular, applying the predicted bounds from the Circular Electron-Positron Collider (CEPC)  CEPC1 ; CEPC2 ; CEPC3 (with 0.01 accuracy for top and bottom quarks and 0.001 accuracy for Z bosons), both the lower middle and right panels would be entirely excluded, while only a minimal portion of the parameter space would remain viable for the upper right panel.
EW scale
$\lambda_{1}$ 0.0335   0.0320 $\lambda$ 0.1340   0.1278 $\sigma_{4}$ 0.7862   0.0417
$\lambda_{2}$0.9263   -0.2818 $\rho_{1}$ 0.3734   -3.1750 $\mu^{\prime}_{1}$ 44.3818   30.3390
$\lambda_{3}$ 0.6798   -0.3413 $\rho_{2}$2.7191   1.3651 $\mu^{\prime}_{2}$ 31.3827   21.4529
$\lambda_{4}$ 0.2608   0.2597 $\rho_{3}$0.4931   -1.2463 $\mu^{\prime}_{3}$3937.8724   -1210.3213
$\lambda_{5}$0.5559   -0.0295 $\rho_{4}$1.8526   -0.5637 $m_{\phi}$ 58.8545   83.2217
$m_{h}$ 126.8221   128.2429 $\rho_{5}$ 2.7191   -1.3651 $m_{\xi}$ 343.1990   113.1068
$m_{H}$ 121.9700   120.6273 $\sigma_{1}$ 0.4871   1.0092 $m_{\chi}$ 355.0436   111.2011
$\sin(\theta_{H})$ 0.5069   0.2067 $\sigma_{2}$ 1.1112   0.0589 $\upsilon_{\Delta}$ 15.1016   17.9776
$\alpha$ 0.1736   1.9891 $\sigma_{3}$ 0.5215   0.5193 $\upsilon_{\phi}$ 243.3934   242.1265
Observeables
$m_{h_{1}}~(m_{H})$ 95.5927   96.1130 $\kappa^{2}_{Hff}$ 0.3846   0.1947 $\kappa_{b}$ 0.9398   1.0034
$m_{h_{2}}~(m_{h})$ 125.5482   125.3452 $BR(H\to\gamma\gamma)$ 0.0009   0.0017 $\kappa_{\tau}$ 0.9398   1.0034
$m_{h_{3}}$ 368.3684   253.7774 $BR(H\to b\bar{b})$ 0.8004   0.7980 $\kappa_{ww}$ 1.0185   1.0247
$m_{a_{2}}$ 310.0807   185.8671 $BR(H\to\tau\bar{\tau})$ 0.0887   0.0885 $\kappa_{zz}$ 1.0193   1.0249
$m_{H^{\pm}_{2}}$ 237.4317   186.1513 $\mu_{\gamma\gamma}$ 0.2490   0.2381 S 0.0330   -0.0795
$m_{H^{\pm}_{3}}$ 323.3204   251.3946 $\mu_{b\bar{b}}$ 0.1115   0.1436 T -0.1027   -0.0164
$m_{H^{\pm\pm}}$ 324.9837   252.1293 $\mu_{\tau\bar{\tau}}$ 0.1190   0.1533 U -0.0200   0.0005
$\kappa^{2}_{HVV}$ 0.1116   0.1441 $\kappa_{t}$ 0.9398   1.0034 P-value 0.8125   0.8100
Observeables
$M_{scale}$ $3.1623\times{10}^{6}~~~6.3096\times{10}^{16}$ HiggsSignals $\checkmark$~~$\checkmark$ ✓
$Br(B_{s}\rightarrow X_{s}\gamma)$ $3.1086\times{10}^{-4}~~~3.0901\times{10}^{-4}$ Unitarity $\checkmark$~~$\checkmark$ ✓
$Br(B^{0}_{s}\rightarrow\mu^{+}\mu^{-})$ $2.8925\times{10}^{-9}~~~2.7844\times{10}^{-9}$ VS at EW ×$\times$~~$\times$ ×
$Br(B\rightarrow\tau\nu)$ $1.1725\times{10}^{-4}~~~1.1431\times{10}^{-4}$ VS at High scale $\checkmark$~~$\checkmark$ ✓
HiggsBounds $\checkmark$~~~~~~~~~~~~~~~~~$\checkmark$ ✓ Positive Definiteness $\checkmark$~~$\checkmark$ ✓
Table 1: Two benchmark points. The first sub-table shows the input parameters in (3) and the values of corresponding parameter in the scalar potential (78) at the EW scale. We also present some low energy experimental observables in the second and the third sub-tables. All dimensional parameters listed in the table are given in units of GeV.
We present two benchmark points, which can satisfy both current experimental and theoretical constraints except the bounded-from-below constraints adopted in previous studies while fulfill the positive definiteness constranints in the large field value regions, in Table 1 that can adress these two excesses of di-photon and $b\bar{b}$.
@@ Acknowledgements.
We would like to thank useful discussions with Fei Wang and Yang Zhang. This work was supported by the National Natural Science Foundation of China (Grant Nos. 12447167 and 12275067), by the Joint Fund of Henan Province Science and Technology R&\&&D Program (Grant Nos.225200810030, 225200810092 and 225200810014), by the Startup Research Fund of Henan Academy of Sciences (Grant No. 231820011), by the Basic Research Fund of Henan Academy of Sciences (Grant No. 240620006), by the Graduate Innovation Fund of Henan Academy of Sciences (Grant No. 243320031), by the Science and Technology Innovation Leading Talent Support Program of Henan Province, by the National Key R&\&&D Program of China (Grant No. 2023YFA1606000) and by the Science and Technology Innovation Leading Talent Support Program of Henan Province (Grant No. 254000510039).